When the price of movies decreases from $15 to $10, the number of tickets sold increases from 3000 to 5000. What is the price elasticity of demand using the midpoint formula? Select one: O a. 400 O b. 1.8 O c. 1.25 O d. 0.025 Next page Answers Jump to...
Answer #1

PRICE ELASTICITY OF DEMAND = % CHANGE IN QUANTITY/% CHANGE IN PRICE

USING MID POINT FORMULA WE CAN FIND OUT

% CHANGE IN QUANTITY = Q2-Q1/(Q2+Q1)/2*100

% CHANGE IN PRICE = P2-P1/(P2+P1)/2*100

USING FIGURES GIVEN IN THE ABOVE EXAMPLE

%CHANGE IN QUANTITY = 5000-3000/(5000+3000)/2*100

%CHANGE IN QUANTITY= 50%

% CHANGE IN PRICE = 10-15/(10+15)/2*100

% CHANGE IN PRICE = -40%

PRICE ELASTICITY OF DEMAND = % CHANGE IN QUANTITY/% CHANGE IN DEMAND

PRICE ELASTICITY OF DEMAND = 50/-40

PRICE ELASTICITY OF DEMAND =- 1.25

CORRECT OPTION IS C= 1.25

(PRICE ELASTICITY IS ALWAYS NEGATIVE AS PRICE AND DEMAND IS INVERSELY RELATED,BUT BY CONVENTION ELASTICITY IS POSITIVE THUS ABOVE MATHEMATICALLY ELASTICITY IS -1.25 BUT INTERPRETED AS 1.25)